Lesedi buys a house and make a down payment of 16% of the price of the house. for the remaining amount,she manages to secure a loan at an interest rate of 12.05% per year,compounded monthly for a period of 20 years her monthly her payment is 18556.84

A..The size of the loan ( to the nearest Rand is)

1. 4 453 642

2. 1 680 000

3. 2 167317

4. 1 338320


B.. The down payment is

1. 213 411

2. 320 00

3. 346 771

4. 268 800






1
Expert's answer
2022-03-20T06:41:41-0400

Lesedi's yearly payment= 18556.84*12=222682.08

A=p(1+r/n)nt

222682.08(1+12.05%)20*12

222682.08(1.0100416)240

=2449672.089

So Lesedi's size of loan is 2449672


b. 84%=2449672

16%=?

(16*2449672)/84

=466604.21

The down payment is 466604